# Resource Public Key Infrastructure BGP Border Gateway Protocol

### v1

@@ VI-H Backwards Compatibility
Simply including these objects into existing repository structure does not provide a benefit, as improved RPs still need to download and validate legacy objects to find the improved objects, and legacy RPs need to additionally download all new objects without being able to process them. We thus propose new objects are added into a parallel version of the repository. An RP supporting iRPKI probes for the existence of a Notification.bin file at the same folder URI as the Notification.xml. This updated RRDP file and its corresponding Snapshot / Deltas contain the updated objects. If a Notification.bin is not found (404), the RP proceeds to download and process the legacy Notification.xml file. This design only results in minimal overhead for the RPs, which generally do not get stalled from the failed request. Experiments shows that Fort is the only RP that will stall for 4 seconds if the server returns a 404 error, compared to milliseconds in Routinator and rpki-client. Implementing iRPKI support into Fort must ensure that a poll for Notification.bin does not result in stalling and the RP should instead immediately move to downloading the legacy Notification.xml. This is possible by setting a smaller request timeout for requests to the Notification.bin.
@@ VII Evaluations
We implement all improvements derived in Section VI into RPKI software to illustrate the performance benefits. We use Routinator 0.14.1 in our evaluations and implement all iRPKI changes to be backwards-compatible: The improved version of Routinator remains fully functional for current RPKI, and additionally supports new object types and validation concepts. In total, adding support for all iRPKI functionality and objects required adding 1836 lines of code to Routinator. To create iRPKI objects and construct them into a repository, we use the open-source RPKI tool CURE [13]. We adapt CURE to support legacy and improved versions of each object, which can be selected through CLI arguments.
@@ VII-A Setup
As RPKI is deployed in production systems, testing on real-world RPKI is unfeasible. We thus instead construct a local isolated environment to run all our evaluations that is not connected to real-world RPKI. All tests are run on an Ubuntu 24.04LTS machine with 32 GB RAM, and an 8-core current generation i7 processor. Nginx is used to serve the RRDP files. Following Section VI, we do not limit the bandwidth of the nginx webserver, as real-world available bandwidth can vary between servers, and size comparison between RPKI and iRPKI provides a clearer picture on bandwidth requirements.
@@ VII-B Repository Structure
Evaluating the impact of individual improvements is not trivial, as benefits depend on the structure of repositories: A CA with many ROAs will benefit more from ROA improvements than from manifest improvements, while a CA without any ROAs will not benefit from ROA improvements at all. Experiments show that real-world CAs are diverse, with a minimal ROA amount of 0, and a maximum ROA amount of 48886. To illustrate benefits for CAs with many ROAs vs. CAs with few ROA, we run each evaluation test-case with two repository structures: To test processing of payloads, like ROAs, we create a single CA setup that includes 10,000 ROA files. We chose 10,000 ROAs as we did not find any changes in our comparative results with more than 10,000 ROAs.
@@
Real-world evaluation. Running our evaluations in a isolated lab environment is essential to provide clear and repeatable impact comparison of the different optimizations, and to not cause issues in real-world RPKI operation. To still evaluate real-world impact of our improvements, we additionally conduct experiments with a setup emulated after the real-world RPKI structure. For this, we measure the amount and structure of all CAs that are currently part of the RPKI, and re-create their structure in our local environment, including 45k CA certificates, manifest, CRLs, and 290k ROAs. To run the experiments locally, we create our own TAL file pointing to our own local domain. Since we do not have access to the root private keys of RIRs, we use our own private root keys. This requires us to re-sign all RPKI objects. Further, we restore the validity of all objects for each test-case. This is necessary, as objects might otherwise expire over the course of the experiments, potentially skewing results.
@@
Total processing time. We time the execution of the RP to evaluate how our improvements impact the total processing time. Performance benefits can stem from multiple parts of the processing, including smaller download times due to smaller objects, faster parsing in internal processing due to smaller objects or improvement object structures, faster validation due to lower cryptographic overhead, and lower cache storage times. We measure execution time with the Linux tool time.
@@ VII-D Results
Flamegraph iRPKI. We run a flamegraph evaluation using the same parameters as in Section III on a full iRPKI setup. Results are shown in Figure 5. In the optimized version, certificate processing takes up most processing time. This is expected, as certificates were intentionally not adapted in the new design to ensure they remain compatible with existing X.509 implementations, and no potential issues with certificate path security are introduced. Manifest processing still requires a signature validation, leading to a remaining large fraction of processing. RRDP processing requires some parsing time, as the Snapshot remains relatively large, containing all repository objects. With the optimizations, processing ROAs is merely reduced to parsing the small objects and storing the content in the local state. In iRPKI, processing ROAs only takes up a minor fraction of total processing time, while it took up almost 50% of processing in legacy RPKI.
@@ VII-E Real-World Evaluations
We additionally evaluate iRPKI on a setup derived from real-world RPKI structure. For this, we measure the structure of RPKI. On the day of our measurement, Jan 28 2025, the global RPKI contains 427,937 objects. Figure 6 illustrates how each type impacts the total (binary) size and total signature validations of current RPKI objects. The figures show that ROAs make up the largest part of all RPKI data and of all required signature validations.
To evaluate how iRPKI would impact real-world objects, we process the current RPKI state to convert all objects and files to their iRPKI counterparts. All ROAs are stripped of their EE-certificate and signature, and serial-number and expiration time are added to the content. The EE-certificate is removed from the manifest, and the content of the respective CRL is added to the manifest. Manifests and ROAs are encoded with protobuf. Notification and Snapshots are adapted to the new improved structure and also encoded with protobuf.
@@
We then run a test with the full current RPKI state converted to iRPKI. With the improved version, Routinator finishes in 57s, 5.8x faster than current RPKI. The total RRDP size is reduced from 1.1 GB to 160 MB, a 6.9x improvement. Since iRPKI does not make changes to object validity, both states lead to the same VRPs output. The real-world evaluations illustrate the substantial benefit of iRPKI when the full current RPKI state is converted to iRPKI objects.
@@ VIII Path Forward: From Design to Deployment
Timeline. Substantial changes to Internet protocols are traditionally slow, and can take years to incorporate into the specification. For example, incorporating ASPA into RPKI has been ongoing for seven years. To facilitate a swift inclusion of our improvements into the specification, we propose starting with easier-to-implement changes first. For example, discussions on a new RRDP version are already ongoing [^12] Link IETF, and could incorporate the changes to the file structure and the encoding proposed in this work. Changes to the fundamental structure and encoding of RPKI objects are essential but require modifying the current object structure and validation specification, and will therefore incur a longer process.
Deployment overhead. During the transition from the current RPKI to iRPKI, both technologies will co-exist. This creates additional load for publication points that support both version: The servers will need to provide all existing objects for all legacy RPs, and additional new version for iRPKI-supporting RPs. However, since iRPKI objects are significantly smaller and require fewer signature computations, the storage overhead is, on average, about 14.5%. Further, the servers get the measurable benefit of less bandwidth and correspondingly less load from all RPs utilizing iRPKI. To facilitate deployments in test-environments, we make exemplary iRPKI repository configurations public for use by operators.
@@ IX Conclusion
To overcome these limitations, fundamental changes to the architecture of RPKI are necessary. In this work, we identify substantial inefficiencies in the design of RPKI and develop specification-level improvements to overcome them. We demonstrate a 6.9x size improvement and 5.8x faster processing when the current real-world RPKI state is converted to iRPKI. iRPKI is backwards compatible with RPKI deployments, and offers benefits even to early adopters. We expect that these properties will contribute to speeding up the required adaptions to the specification and the standardization process. With iRPKI, the RPKI architectures becomes scalable, enabling support for new and future use-cases, operation in constrained environments, and promises to remain functional in full global deployment. To facilitate deployment and changes to specification and enable reproduction of our study, we open-source exemplary iRPKI objects, the adapted Routinator code, and the repository tool [^13] Study Data.

### v2

@@ VI-H Backwards Compatibility
Simply including these objects into existing repository structure does not provide a benefit, as improved RPs still need to download and validate legacy objects to find the improved objects, and legacy RPs need to additionally download all new objects without being able to process them. We thus propose new objects are added into a parallel version of the repository. An RP supporting iRPKI requests a Notification.bin file at the same folder URI as the Notification.xml, e.g., ../rrdp/notification.bin instead of ../rrdp/notification.xml. This updated RRDP file and its corresponding Snapshot / Deltas contain all updated objects. If a Notification.bin is not found (404), the RP falls back to regular RPKI to download and process the legacy Notification.xml file. This design only results in minimal overhead for the RPs, which generally do not get stalled from the failed request. Experiments shows that Fort is the only RP that will stall for 4 seconds if the server returns a 404 error, compared to milliseconds in Routinator and rpki-client. Implementing iRPKI support into Fort must ensure that a poll for Notification.bin does not result in stalling and the RP should instead immediately move to downloading the legacy Notification.xml. This is possible by setting a smaller request timeout for requests to the Notification.bin.
@@ VII Evaluations
We implement all improvements derived in Section VI into Routinator 0.14.1 to illustrate the performance benefits. All iRPKI changes are backwards-compatible: The improved version of Routinator remains fully functional for current RPKI, and additionally supports new object types and validation concepts. In total, adding support for all iRPKI functionality and objects required adding 1836 lines of code to Routinator. To create iRPKI objects and construct them into a repository, we use the open-source RPKI tool CURE [25]. We adapt CURE to support legacy and improved versions of each object, which can be selected through CLI arguments.
@@ VII-A Setup
As RPKI is deployed in production systems, testing on real-world RPKI is infeasible. We instead construct a local isolated environment to run all our evaluations that is not connected to real-world RPKI. All tests use Ubuntu 24.04LTS with 32 GB RAM, and an 8-core current generation i7 processor. Nginx is used to serve the RRDP files. Following Section VI, we do not limit the bandwidth of the nginx webserver, as real-world available bandwidth can vary between servers, and size comparison between RPKI and iRPKI provides a clearer picture on bandwidth requirements.
@@ VII-B Repository Structure
Evaluating the impact of individual improvements is not trivial, as benefits depend on the structure of repositories: A CA with many ROAs will benefit more from ROA improvements than from manifest improvements, while a CA without any ROAs will not benefit from ROA improvements at all. Experiments show that real-world CAs are diverse, with a minimal ROA amount of 0, and a maximum of 48886. To illustrate benefits for CAs with many ROAs vs. CAs with few ROA, we run each evaluation test-case with two repository structures: To test processing of payloads, like ROAs, we create a single CA setup that includes 10,000 ROA files. We chose 10,000 ROAs as we did not find any changes in our comparative results with more than 10,000 ROAs.
@@
Real-world evaluation. Running our evaluations in a isolated environment is essential to provide clear and repeatable comparisons of the different optimizations, and to not cause issues in real-world RPKI operation. To still evaluate real-world impact of our improvements, we additionally conduct experiments with an emulated real-world RPKI structure. For this, we measure the amount and structure of all CAs that are currently part of the RPKI, and re-create their structure in our local environment, including 45k CA certificates, manifest, CRLs, and 290k ROAs. To run the experiments locally, we create our own TAL file pointing to our own local domain. Since we do not have access to the root private keys of RIRs, we use our own private root keys. This requires us to re-sign all RPKI objects. Further, we restore the validity of all objects for each test-case. This is necessary, as objects might otherwise expire during the experiments, potentially skewing results.
@@
Total processing time. We time the execution of the RP to evaluate how our improvements impact the total processing time. Performance benefits can stem from multiple parts of the processing, including smaller download times and parsing due to smaller objects or improved structure, faster validation due to lower cryptographic overhead, and lower cache storage times. We measure execution time with the Linux tool time.
@@ VII-D Results
Flamegraph iRPKI. We run a flamegraph evaluation using the same parameters as in Section III on a full iRPKI setup, shown in Figure 5. In the optimized version, certificate processing takes up most processing time. This is expected, as certificates were intentionally not adapted in iRPKI to ensure compatibility with existing X.509 implementations, and ensure no potential issues with certificate path security are introduced. Manifest processing still requires a signature validation, leading to a remaining large fraction of processing. RRDP processing requires some parsing time, as the Snapshot remains relatively large, containing all repository objects. With the optimizations, processing ROAs is merely reduced to parsing the small objects and storing the content in the local state. In iRPKI, processing ROAs takes up minimal processing time, while it took up 5̃0% of processing in legacy RPKI.
@@ VII-E Real-World Evaluations
We additionally evaluate iRPKI on a setup derived from real-world RPKI structure. On the day of our measurement, Jan 28 2025, the global RPKI contains 427,937 objects. Figure 6 illustrates how each type impacts the total (binary) size and total signature validations of current RPKI objects. The figures show that ROAs make up the largest part of all RPKI data and of all required signature validations. A comparison to iRPKI can be found in Appendix -F.
To evaluate how iRPKI would impact real-world objects, we process the current RPKI state to convert all objects to their iRPKI counterparts. All ROAs are stripped of their EE-certificate and signature, and serial-number and expiration time are added to the content. The EE-certificate is removed from the manifest, and the content of the respective CRL is added to the manifest. Manifests and ROAs are encoded with protobuf. Notification and Snapshots are adapted to the new improved structure and also encoded with protobuf.
@@
We then run a test with the full current RPKI state converted to iRPKI. With the improved version, Routinator finishes in 57s, 5.8x faster than current RPKI. The total RRDP snapshot size is reduced from 1.1 GB to 160 MB, a 6.9x improvement. We additionally include delta bandwidth evaluation in Appendix -E. Since iRPKI does not make changes to object validity, both states lead to the same VRPs output. The real-world evaluations illustrate the substantial benefit of iRPKI when the full current RPKI state is converted to iRPKI.
@@ VIII Path Forward: From Design to Deployment
Timeline. Substantial changes to Internet protocols are traditionally slow, and can take years to incorporate into the specification. For example, incorporating ASPA into RPKI has been ongoing for seven years. To facilitate a swift inclusion of our improvements into the specification, we propose starting with easier-to-implement changes first. For example, discussions on a new RRDP version are already ongoing [7], and could incorporate the changes to the file structure and the encoding proposed in this work. Changes to the fundamental structure and encoding of RPKI objects are essential but require modifying the current object structure and validation specification, and will therefore incur a longer process.
Hybrid deployment. During the transition from RPKI to iRPKI, both technologies will co-exist. This creates additional load for repositories supporting both versions: Repositories must retain all existing objects for legacy RPs, and offer iRPKI objects for updated RPs; a storage overhead of about 14.5%. Supporting iRPKI allows the repository to serve significantly smaller snapshots (6.9x on average) to any RP supporting iRPKI. Following Section V, bandwidth is a larger operational concern for repositories than storage and iRPKI thus provides operational benefits even during hybrid deployment. For RPs, the only overhead is potential failed requests to Notification.bin files for all repositories not supporting iRPKI, which currently amounts to a maximum of 91 requests. RPs supporting iRPKI get the bandwidth, computation and storage improvements for any iRPKI-supporting repository.
@@ IX Conclusion
To overcome these limitations, fundamental changes to the architecture of RPKI are necessary. In this work, we identify substantial inefficiencies in the design of RPKI and develop specification-level improvements to overcome them. We demonstrate a 6.9x size improvement and 5.8x faster processing when the current real-world RPKI state is converted to iRPKI. iRPKI is backwards compatible with RPKI deployments, and offers benefits even to early adopters. We expect that these properties will contribute to speeding up the required adaptions to the specification and the standardization process. With iRPKI, the RPKI architectures becomes scalable, enabling support for new and future use-cases, operation in constrained environments, and promises to remain functional in full global deployment. To facilitate deployment and changes to specification and enable reproduction of our study, we open-source exemplary iRPKI objects, the adapted Routinator code, and the repository tool666www.dropbox.com/scl/fi/pvs6odh2ml25azifd6vg9/submission_ndss.zip.